### v2

# A practical guide to Digital Micro-mirror Devices (DMDs) for wavefront shaping

Abstract
Digital micromirror devices have gained popularity in wavefront shaping, offering a high frame rate alternative to liquid crystal spatial light modulators. They are relatively inexpensive, offer high resolution, are easy to operate, and a single device can be used in a broad optical bandwidth. However, some technical drawbacks must be considered to achieve optimal performance. These issues, often undocumented by manufacturers, mostly stem from the device’s original design for video projection applications. Herein, we present a guide to characterize and mitigate these effects. Our focus is on providing simple and practical solutions that can be easily incorporated into a typical wavefront shaping setup.
September 2023
1 Introduction
Since the advent of adaptive optics, various technologies have been employed
to modulate the amplitude and/or phase of light.
Early adaptive optics devices, utilized in fields like microscopy and astronomy,
offer rapid modulation capable of compensating for the aberrations of optical systems
in real-time.
However, these devices are constrained by a limited number of actuators,
restricting their utility in complex media where a large number of degrees of freedom is essential.
Liquid Crystal Spatial Light Modulators (LC-SLMs),
which allow for the control of light phase across typically more than a million pixels,
have emerged as powerful tools for wavefront shaping in complex media
since the seminal work of A. Most and I. Vellekoop in the mid-2000s [1].
Nonetheless, LC-SLMs are hampered by their slow response time,
permitting only a modulation speed ranging from a few Hz to a few hundred Hz.
Digital Micromirror Devices (DMDs) have emerged as a technology bridging the gap
between these two types of systems;
they offer a large number of pixels (similar to LC-SLMs) and fast modulation speeds (typically up to several tens of kHz).
Their high speed capabilities made them attractive for real-time applications,
in particular for high-resolution imaging microscopy
requiring fast scanning or illumination shaping [2, 3],
biolithography [4],
and optical tweezers [5].
However, DMDs are restricted to hardware binary amplitude modulation and are not optimized for coherent light applications.
Utilizing DMDs for coherent control of light in complex media is therefore non-trivial
and necessitates specific adaptations for efficient use.
To comprehend both the capabilities and limitations
of DMD technology for coherent wavefront shaping,
it is crucial to understand the device’s operating principles
and its original design intentions.
Investigated and developed by Texas Instruments since the 1980s,
DMDs gained prominence in the 1990s for video projection applications since the 1990s
under the commercial name of Digital Light Processing (DLP) [6, 7].
The technology enables high-resolution, high-speed, and high-contrast-ratio modulation of light.
DMDs operate by toggling the state of small mirrors between two distinct angles, denoted as .
The device is originally engineered for amplitude modulation in video projection applications.
In this configuration, one mirror angle directs light into the projection lens,
while the alternate angle results in the light path being blocked (see Fig. 1).
Given that projectors utilize incoherent light and that the DMD plane is optically conjugated with the projection screen,
aberrations within the DMD plane are generally not problematic.
Similarly, phase fluctuations induced by temperature variations,
as well as minor vibrations from the cooling hardware, are inconsequential in this context.
The DMD is designed to produce binary on/off modulation,
which is then leveraged to generate grayscale images via pulse-width modulation.
Color modulation is accomplished through the use of a color wheel in conjunction with a bright white light source.
Third-party companies have developed kits tailored for research applications,
which include a DMD, a control board, and a software interface.
Specifically, Vialux devices [8] offer an FPGA board that enables high-speed modulation
by allowing frames to be stored in the device’s memory [9].
However, standard Texas Instruments video projector evaluation modules can also be repurposed into wavefront shaping devices [10],
though at a compromised modulation speed.
These systems can further be converted into phase or complex field modulators.
This is typically achieved by encoding the optical phase into
the spatial displacement of binary spatial fringes displayed on the DMD,
followed by filtering the high spatial frequencies in the Fourier plane [11].
Such a configuration permits multi-level complex modulation but sacrifices spatial resolution.
The implementation and performance of such systems are explored in a separate tutorial [12]
and will not be elaborated upon here.
For the remainder of this paper, it will be assumed that the DMD is used for complex modulation via such a method.
While other articles exist describing the various aspects of DMDs [13, 14, 10, 15],
this tutorial aims to provide a guide for easily setting up a DMD for wavefront shaping applications in complex media.
In particular, we provide characterization and validation procedures that require minimal changes compared to typical wavefront shaping setups.
More specifically, we place ourselves in a standard experiment where the goal is to shape the complex wavefront impinging on a complex medium and control or measure its output response.
This is typically the case for a focusing experiment [1] or for measuring the transmission matrix of a complex medium [16].
We first introduce the diffraction properties of a DMD
and elaborate on how these could impact the system’s efficiency.
We also furnish a straightforward criterion for selecting the appropriate DMD parameters for a specified excitation wavelength.
In the next section,
we delve into the aberration impacts brought about by the non-flatness of the DMD surface.
We demonstrate a simple process to characterize this effect and provide compensation solutions.
In the third segment,
we detail the influence of mechanical vibrations that are induced by the DMD’s cooling system.
Lastly, we discuss how the thermalization of the DMD chip can potentially result in variations to the DMD response over time.
2 Choosing the right DMD: Diffraction effects
2.1 A 1D model
A significant distinction between liquid crystal modulators and DMDs
lies in the geometry of the pixel surface.
This difference gives rise to diffraction effects that can adversely affect
both the modulation quality and system efficiency.
The impact of these diffraction effects is highly contingent on several factors:
the wavelength of the illumination, the pixel pitch, and both the incident and outgoing angles.
Therefore, alongside selecting an appropriate anti-reflection coating,
it is crucial to ensure that the pixel pitch is compatible with the specific configuration being used.
Texas Instruments offers chips with a variety of pixel pitches ,
ranging approximately from to µm [18].
To achieve a qualitative understanding of this issue,
we consider a 1D array of pixels as illustrated in Fig. 2.
Initially, let’s assume that all pixels are in the same state and are illuminated by a plane wave originating from the far field.
Under these conditions, the pixelated modulator essentially functions as a grating,
with a period that is equivalent to the pixel pitch.
It is important to underscore that these modulators possess a hardware-limited fill factor,
typically around 90%.
This translates to an effective active pixel size of .
In general, a grating gives rise to various diffraction orders
with differing intensities
and angles , as dictated by the grating equation:
where is the wavelength of the light,
is the angle of the first-order diffraction,
and is an integer value denoting the orders of diffraction.
The intensity of the individual diffraction orders is influenced by
the response of a single pixel,
constituting the unit cell of the grating,
and that is governed by , , and its rotation angle in the case of a DMD.
Importantly, we can decouple the effects of the grating’s periodicity,
which influences the angles of the diffraction orders,
from the effects of the response of a single pixel,
which shapes the envelope of the angular response.
For a case of normal incidence with an LC-SLM (Liquid Crystal Spatial Light Modulator),
we can assume that the response of a single pixel is uniform over its surface.
Similarly, in the scenario of a blazed grating,
such as a DMD, a linear phase slope is present on each pixel.
This is due to the tilt angle of the mirrors.
For an arbitrarily selected incidence angle ,
a global phase slope is introduced.
This results in a trivial shift of the angular diffraction pattern by an angle .
In essence, the incident angle serves to shift the entire angular diffraction pattern
relative to the case of normal incidence,
while the blazed angle
—the tilt angle of the mirror in the DMD projected onto the axis we consider—
only shifts the envelope by an angle of .
Whenever the fill factor approaches 100%,
i.e. when ,
the envelope for a flat grating achieves its maximum at ;
this corresponds to the angle of the zeroth-order diffraction,
and the intensity nears zero for all other orders.
In these specific conditions, a singular diffraction order is visually perceived, corresponding to the optimum scenario.
The addition of a blazed angle results in both
a shift in the envelope and in the position of its maximum,
now indicated by .
In the general case,
this position may not align with a diffraction order anymore [13].
We provide in Appendix A. a 1D computation of this effect.
A more accurate computation of the far field can be found in [15].
We represent in Fig. 3 the angular response of a flat grating and a blazed grating for a 1D filling fraction of 95% (correponding to a 2D filling fraction of %). For a flat grating, the zero-th order contains most of the intensity, the other orders being negligible in comparison. For the blazed grating example shown, we are in a situation close to the worst case scenario: Two diffraction orders have a significant and comparable intensity, and other orders also have a non-negligible contributions. In the optimal scenario, where the peak of the envelope corresponds to a diffraction order, it results in a single diffraction order carrying the majority of the energy. This state is achieved when the conditions of the blazed grating equation are fulfilled [19]:
We note that the incident angle olso affects the diffraction efficiency.
2.2 The 2D case
To analyze more precisely the effect of diffraction in a DMD,
one needs to consider the 2D surface of the modulator.
We can establish a Cartesian coordinate system on the plane of the DMD,
with axes and aligned with the pixel sides
(refer to Fig. 4a).
Pixels are uniformly repeated along these axes.
However, a technical challenge arises
in that the axis of rotation of the pixels
aligns with the pixel diagonals,
resulting in a rotation by 45 degrees with respect to the and axes.
For the convenience of alignment and manipulation of the optical setup,
it is preferable to work with the incident and outgoing beams
which have the optical axis contained in the horizontal plane,
i.e. a plane parallel to the table surface.
A straightforward and prevalent solution is to rotate the chip by 45 degrees
relative to the horizontal plane,
which aligns the pixel’s axis of rotation to be vertical.
This configuration is depicted in Fig. 4b.
The 2D array can be seen as 2 orthogonal 1D gratings in the and directions, both having the same properties. A more comprehensive depiction of the 2D system can be found in [14]. In order to utilize Eq.(1), one first needs to project the different angles of the problem onto the incident planes of the two 1D gratings. This yields , , and , where represents the true rotation angle of the mirrors relative to the diagonal of the pixels, and and denote the incident and outgoing angles in the horizontal plane. We can quantify how close we are to the ideal case, i.e., when satisfying the blazed equation outlined by Eq.(1), by defining a blazed number as introduced in [20]:
with representing the floor function,
and the modulo 2 operation.
is maximal and equals when the blazing equation is satisfied,
i.e. when one order of diffraction contains most of the energy,
and minimal when we are in the worst-case scenario,
i.e. when four orders of diffraction have a significant and equal intensity.
To demonstrate the effect, we conduct a simulation of a DMD using Python (refer to tutorial and code in [20]) with two pixel pitches of µm and µm, under a coherent excitation at nm. Fig. 5 shows the estimated blazed number as a function of the angle of incidence in the horizontal plane, along with the far field diffraction pattern for two distinct incident angles. Fig. 6 shows the blazed number as a function of both the incident angle and the wavelength for pixel pitches of µm and µm. It should be noted that the efficiency of diffraction can be altered by adjusting the angle of incidence. However, its impact is relatively confined within an acceptable angular range that aligns with experimental limitations (i.e. for angles far from ). Far-field patterns are centered around the maximum of the envelope (marked by a yellow cross). We see that for small values of , the pixel pitch of µm leads to a blazed number close . It corresponds in the far field to having one bright order of diffraction close to the maximum of the envelope.
2.3 Modulation cross-talk
In practice, we place a pinhole or iris to select
one order of diffraction,
corresponding to the on state.
Having a small value for the blaze number
not only restricts the amount of light modulation
due to the diminished diffraction efficiency,
it also influences the modulation quality by inducing cross-talk
between the two states of the DMD pixels.
Until now, we have assumed that all the pixels are in the same state.
In actual usage of the DMD,
it becomes necessary to modulate the state of each pixel individually.
When approaches zero, higher orders of diffraction
still possess a significant intensity,
as demonstrated in the 1D case in Fig 3.
One adverse implication is that pixels in the off state
may contain orders of diffraction that are not blocked by the pinhole,
and therefore, will contribute as an interference to the modulated wavefront.
We show in Fig. 7 the normalized amplitude of the diffraction patterns
corresponding to all the pixels in the on (blue curve) and off (orange curve) states
for the same experimental conditions but with two different pixel pitches
leading to situations close to the worst (a) and best (b) case scenarios.
We observe the presence of a non-negligible contribution of the off state
at the main diffraction order of the on state in the first case.
While this contribution might appear weak,
it does affect the quality of the modulation
since the modulation scheme typically necessitates about half the pixels
to be in the off state for phase modulation [11],
and even more so for elaborate modulation schemes [21, 12].
2.4 Dispersion
DMDs are composed of metallic small mirrors,
the response of which is minimally affected by wavelength changes.
This is particularly advantageous for broadband applications requiring amplitude modulation
and operating on a plane conjugated to the DMD’s surface.
This is the case for the originally intended application of video projection.
However, for wavefront-shaping applications,
it is typically required to select a specific diffraction order
to acheive phase or complex modulation [11, 12].
Under such circumstances,
the wavelength-dependency of the diffraction effect becomes important.
The blazed number, denoted by (according to Eq.(2)),
scales inversely with the wavelength.
Fig. 8 shows the blazed number
as a function of the wavelength for two pixel pitches µm and µm,
with an incident angle of .
Within the visible spectrum, fluctuates between to over a typical range of roughly 100 nm.
2.5 Python code example
We provide in the paper repository [23] a Python code to simulate the diffraction effect of a DMD by computing the far field pattern for a set of realistic parameters. It provides a simple way to estimate the blazed number introduced in eq. (2) to assess the quality of the modulation at the desired wavelength for a given pixel pitch. We also propose an online tool accessible at https://www.wavefrontshaping.net/post/id/49.
3 Characterizing aberration effects
3.1 Presentation of the problem
While only capable of providing hardware binary amplitude modulation,
DMDs serve as a potent tool for wavefront shaping and sensing.
These applications critically require characterization and correction of aberrations
caused by the non-flatness of the DMD surface.
For LC-SLMs, the manufacturer typically characterizes the surface’s inhomogeneities within the plane of the modulator and
provides a spatial phase profile of the introduced aberrations based on the operational wavelength.
It is noteworthy that when utilized for intensity modulation on a plane conjugated to the DMD plane,
as it is done in digital projectors,
the system becomes insensitive to the aberrations caused by the DMD surface.
Consequently, these effects are commonly overlooked and rarely documented in the information provided by the manufacturers.
Nonetheless, they are non-negligible, with deviations from a flat surface
that typically spans several wavelengths [24].
3.2 Finding the correction pattern
In the literature, various methods have been proposed to characterize
the phase pattern of the DMD aberrations in the plane of the modulator.
Typically, this involves using a model for the aberrations,
tweaking the parameters to align with the measurements [25, 14, 24].
An alternative approach entails direct measurement of the distorted wavefront,
either employing a wavefront sensor such as a Shack-Hartmann [26],
or via interferometry.
While such methods yield accurate results,
they necessitate adaptation to the particular
setup conditions,
and frequently require supplementary optical components,
meticulous alignment,
and custom software.
Furthermore, they also demand
a meticulously calibrated and stable setup.
In this section, we introduce a straightforward method to characterize the aberrations using a lens and a camera.
This technique can be employed for any system that offers phase modulation,
such as LC-SLMs or deformable mirrors.
We assume the DMD is configured to deliver phase modulation [11, 12].
This implies that the modulator can be divided into sections,
which we designate as macropixels,
where the phase can be controlled independently.
We use a lens and a camera in its Fourier plane,
illuminating the modulator with a collimated beam
that extends over the entire area of the modulator intended for use.
In scenarios where there are minimal or no aberrations,
the intensity pattern observed would mimic the PSF of the lens,
such as an Airy disk depicted in Fig. 9.c.
However, in practice, we encounter a substantially distorted pattern,
like the one represented in Fig. 9.d.
A more detailed depiction of the setup for aberration characterization,
in the context of a wavefront shaping application in complex media is presented in Fig. 10.a.
In the case where the medium is a multimode fiber,
one could leverage the reflection from the input surface
to directly visualize the intensity pattern of the input plane,
as shown in Fig. 10.b.
Note that for the later approach to be reliable, it is necessary to properly align the system to ensure that the plane of the input facet corresponds exactly to the Fourier plane of the DMD, and that the input facet plane is conjugated with the camera plane.
We hypothesize that the aberrations brought about by the DMD are smooth, and can be depicted by a phase pattern in the plane of the DMD array. This could be feasibly approximated by a finite number of Zernike polynomials [27], as follows:
The goal is to find and display the phase value on each macropixel that best compensates for the aberrations, i.e. . We create this pattern in the basis of Zernike polynomials
the best correction is obtained for
We perform a sequential optimization of parameters
to maximize a specific function designed
to be maximal for the ideal correction of optical aberrations.
We first generate a mask, represented by a disk,
centered around the point of maximum intensity of the original image (refer to Fig. 9.d)
with a radius equivalent to a single speckle grain.
The exact size of this radius for a successful optimization is not critical
and can be determined by approximating the dimensions of the ideal point spread function,
expressed as ,
where is the numerical aperture of the optical system and refers to its magnification.
For a given output intensity pattern, we compute an element-wise product between this image
and the created mask, followed by a summation.
This calculated sum is then divided by an analogous product, but with the complementary mask substituted in place of the original
in order to minimize the side lobes of the point spread function.
We set initial parameter values as .
For each parameter, we test different values of ,
construct the phases for every micropixel according to
record the resulting intensity profile,
and evaluate the corresponding cost function.
For each parameter, the value that results in the maximum output is retained.
To mitigate potential noise or instability, this complete process is reiterated times for each parameter.
To estimate the quality of the correction, we compute the Strehl ratio of the point spread function. It is defined as the maximum of the measured PSF divided by the maximum of the ideal one. This optimal PSF is the squared modulus of the Fourier transform of a circular aperture [28]:
with the Bessel function of the first kind of order ,
the wavenumber,
the radius of the aperture,
the radial coordinate in the Fourier plane,
and the focal length of the lens.
As an illustration, we conduct an optimization procedure using 11 Zernike polynomials.
We use a V-9501 Vialux DMD with a DLP9500 TI chip
of resolution 1920 by 1200 pixels and a pixel pitch of 10.8 µm.
The optimization is performed on a disk of radius pixels,
The illumination is done using an expanded laser beam at 633 nm,
corresponding to the aperture of our optical setup.
We exclude the first three Zernike polynomials in the optimization process,
starting from the radial degree .
Indeed, the initial one, known as the piston, does not influence the PSF quality.
The subsequent ones, the tip and the tilt, cause the PSF to shift.
Our procedure relies on optimizing the maximum of the PSF, wherever that is,
rendering us indifferent to these two parameters.
After optimization,
it is possible to generate the correction pattern using a selected number of Zernike polynomials
in order to investigate their impact on the PSF quality.
We see here that using about 10 Zernike polynomials is sufficient to obtain a Strehl ratio .
Fig. 12 demonstrates the Strehl ratio and the intensity profiles of the PSF
for different counts of the utilized Zernike polynomials.
It is important to note that we do not use the full surface of the DMD.
Using a larger area may lead to stronger deformations of the PSF,
requiring a larger number of Zernike polynomials to be corrected accurately.
The experimental data, in addition to the Python code used to generate the figures, can be accessed in the dedicated repository [23].
3.3 Python code example
We provide in the paper repository [23] a Python code example to simulate the effect of aberration on a DMD and then perform a sequential optimization as previously proposed to learn the characterize pattern. We make use of the aotools package [29] for generating Zernike polynomials.
4 Mechanical and thermal stability
Unlike the original purpose of the DMD, i.e. amplitude modulation for video projectors,
typical scientific applications require a high stability of the generated wavefront.
This is particularly true for applications in complex media,
such as strongly scattering media or multimode fiber,
where a small change in the phase front can lead to a large change in the output intensity profile.
While LC-SLMs design has been improved and adapted to scientific applications
over the last decades,
DMD are still relatively new tools for wavefront shaping and sensing
and are prone to instabilities that need to be addressed by the user.
In this section, we present the effect of mechanical and thermal instabilities
and how to limit their impact on the wavefront quality with simple solutions.
4.1 Mechanical stability
Most DMD kits consist of two primary components,
the chip itself and the electronic board that controls it.
This could be the standard electronics board typically used for video projectors,
as seen in TI evaluation kits,
or an FPGA specifically designed for rapid scientific usage,
as offered by Vialux [8] for instance.
Integral to these electronics is a fan designed to cool both the chip and the electronic board.
However, due to the use of a rigid flat cable for connection between the chip and the electronics board,
these parts are not mechanically independent.
As such, vibrations originating from the board are partially transmitted to the chip,
resulting in minute rotations of the mirror surface.
Although this perturbation is inconsequential for video projection,
they can have significant impacts on applications involving complex media
given their high sensitivity to phase front variations.
Due to its high sensitivity in complex media,
it is convenient to characterize this effect
directly on the system’s response,
rather than constructing a distinct setup to analyze the wavefront itself.
An example of such a setup is demonstrated in Fig. 13,
although a similar approach can be employed with a scattering medium.
We enlarge a laser beam onto the DMD and transmit the incoming light through a multimode fiber.
Additional elements are required in the setup to fulfill the requirement for complex modulation [12].
For the sake of clarity, we present a simplified version of the setup where those elements are not present.
The output from the fiber is then made to interfere with a reference arm
in an off-axis configuration [30],
allowing us to detect changes in the output complex field
by recording the interference pattern using a camera.
In the supplementary materials [31], we present an animation illustrating the dynamic pattern.
In off-axis holography,
the local transverse displacement of the fringes
is directly proportional to the phase,
with a displacement equivalent to the period of the fringes corresponding to [30].
This permits us to estimate the fluctuation in phase over time at a given position of the output plane.
As illustrated in Fig. 14 (depicted by the red curve),
the phase varies rapidly over time,
a fluctuation attributed to the rapid mechanical vibrations transmitted by the board.
A simple yet effective solution consists in dampening the vibrations
at the flat cable’s level by clamping it with a soft material,
as depicted in Fig. 15.
This can be achieved using commonly available materials.
In this context, we utilize simple foam, typically used for packaging, and secure it to the cable
with two metallic plates, screws, and nuts.
We observe a significant decrease in the phase fluctuations,
as demonstrated in Fig. 14 (blue curve).
4.2 Thermal stability
Electronics utilized to control the DMD chip experience thermal variations during operation.
The dynamics of this effect are dependent on the frame rate.
Specifically, the chips heats up more quickly when increasing the frame rate.
The increase of temperature can reach more than degrees Celsius
when running a sequence at maximum speed (20-30 kHz) [32].
Notably, this effect is less pronounced when the device is on but not running a sequence.
Temperature fluctuations can cause deformations on the chip’s surface
and can modify the phase response of the glass protective window.
This creates low order aberrations, which degrade the quality of the wavefront.
While this issue is comparatively less critical than static aberrations and mechanical instabilities
previously detailed,
it nonetheless has a substantial impact on the complex medium’s response
when a DMD is employed to modulate the input wavefront.
Empirically, we observe deformations of the pattern
at the output of the medium under study when the DMD chip heats up after switching on the device.
These deformations originate from a temperature-dependent perturbation of the modulated field.
We attribute this effect to the transparent part of the DMD
(i.e., the protection glass window)
having a refractive index that significantly changes with temperature.
Once the device reaches thermal equilibrium, this effect stabilizes,
allowing us to find the appropriate correction mask as described in Section 3 of the manuscript.
It is important to stress that it is the thermal fluctuations that impact the stability of the modulation,
not the temperature itself,
as we obtain stable experimental conditions even when equilibrium is reached
at the maximal operating temperature of the modulator without thermal cooling.
Before initiating a wavefront shaping experiment, it is important to characterize the influence of temperature to assess the extent to which it impacts the results. Although the exact effect on the wavefront distortion can be directly quantified [32], it is typically more convenient to directly measure the effect on the studied system’s response. Directly using the output of the complex medium under study presents multiple advantages. First, it allows the setup to remain unchanged and facilitates easy re-characterization of the effect if some parameters change. Additionally, complex media tend to be very sensitive to input phase changes, with the sensitivity varying from one medium to another. By using the complex medium itself, we obtain the exact sensitivity needed for the current setup. Therefore, regardless of the qualitative changes in the input field, stability in the output indicates that the setup is adequate for studying the medium in question. To do so, we use a setup similar to the one presented in the previous section and depicted in Fig. 13. We can then estimate the field or intensity decorrelation over time. We present here results with intensity correlation, as it does not require a reference arm. The measured output pattern typically take the form of a seemingly random speckle pattern, that is sensitive to minute changes in the input wavefront. The correlation estimation is obtained by comparing the output intensity pattern at a given time to the one at . We use the following expression for the correlation:
with
and .
represent the spatial averaging over the region of interest
of the output plane (i.e. the plane of the camera sensor)
and the temporal averaging over the measured frames.
Fig. 16 shows the measured decorrelation
over time when running an sequence in an infinite loop.
correspond in the first case (a)
to the start of a sequence
and is set in the second one to
4.2 hours post the after of the sequence (b).
For the scenario (a), it is to note that before the sequence commenced,
the DMD was active but remained in the idle state,
meaning that it was not executing a sequence.
We observe a non-negligible decorrelation over time
after the sequence started
that slows down after about 1 hour.
When running the same experiment 4.2 hours after starting the sequence,
the correlation is now stable.
The more efficient way to counteract this effect is to use a closed-loop system to stabilize the temperature of the DMD chip. This can be achieved by using a thermoelectric cooler as demonstrated in Ref. [32] and depicted in Fig. 17.
5 Conclusion
DMDs are powerful tools for wavefront shaping applications in complex media due to their high pixel count, relative low cost, and high refresh rate. However, mostly due to their original purpose of amplitude modulation of incoherent light for video projection, several effects need to be taken into account when using DMDs for wavefront shaping. First, the choice of its pixel pitch must be made carefully by considering the wavelength of operation to ensure a good modulation quality and diffraction efficiency. Furhtermore, the DMD surface is not flat and can introduce aberrations which can be corrected using a simple optimization procedure. Finally, the DMD is sensitive to mechanical vibrations and thermal variations, which can be mitigated ensuring a good mechanical and thermal isolation of the device.
Data and code availability
Data and code examples are available in the dedicated repository [23].
Appendix A: 1D calculation of the DMD diffraction effect
Assuming the effect of the device’s finite size and illumination to be negligible, and situating ourselves within the context of the small-angle approximation, we can represent the field reflected from the device under the influence of plane wave illumination across two systems as follows:
with the rectangular function, representing the finite size of the pixel, defined as:
The intensity as a function of the angle in the far-field is given, up to a homotetic transformation, by the absolute value squared of
The Fourier transform of Eq. 8 can be written as:
We observe that the envelope (right hand term) is maximal for while the effect of the periodicity (left hand term) is maximal for , representing the orders of diffraction.
S.M.P, M.W.M, and R.G.C and acknowledge the French Agence Nationale pour la Recherche grants No. ANR-23-CE42-0010-01 MUFFIN and No. ANR-20-CE24-0016 MUPHTA and the Labex WIFI grant No. ANR-10-LABX-24, ANR-10-IDEX-0001-02 PSL*.